OpenStudy (anonymous):

SIMPLIFY 0.3+ 1 3/4 A. 2 1/2 B. 1 3/7 C. 4 3/4 D. 2 1/20

OpenStudy (anonymous):

\[\frac{3}{10}+\left(1+\frac{3}{4}\right)=\frac{41}{20}=2+\frac{1}{20} \]